Well my teacher started with Slope which I understand perfectly. Than he showed us Standard form, Slope-Intercept form, and Point Slope form. For our homework he gave us 4 problems, 1.The points (-5,2) and (6,-4) 2.m=3/4 and the point (3,5) 3.-3/4x=5/3y-5 and 4.6-4x=3y. With those he wants us to find the equation for all the forms, slope and a graph. I just don't understand how to do that! If you could please explain how, that would be wonderful. Thanks a lot!

Answer by checkley71(8403)

(-5,2)(6,-4) HAS A SLOPE OF (-4-2)/(6+5) OR -6/11
Y=-(6/11)X+b OR SUBSTITUTING -4 FOR X & 2 FOR Y WE GET
2=-(6/11)-4+b OR 2=24/11+b OR 2-24/11=b OR 22/11-24/11=b OR -2/11=b
SO THE LINE EQUATION IS Y=-(6/11)X-2/11